3) Verify the identity:sin sec x cosx x 4) Solve the equation tana
Answer #1

Ans 3

From the double angle formula

Sin2x = 2 Sinx Cosx

Cos2x = 2 Cos2x - 1

Plug these formula into the given question

(2 Sinx Cosx/ Sinx) - [(2 Cos2x - 1)/Cosx] = sec x

(2 Cosx) - [(2Cos2x/Cosx) - (1/Cosx)] = sec x

(2 Cosx) - [(2Cosx) - (sec x)] = sec x

2Cosx - 2Cosx + sec x = sec x

sec x = sec x

Ans 4

tan \theta = \sqrt 3

tan \theta = tan 60°

\theta = 60°